### v1

# SonifyAR: Context-Aware Sound Generation in Augmented Reality

Abstract.
Sound plays a crucial role in enhancing user experience and immersiveness in Augmented Reality (AR). However, current platforms lack support for AR sound authoring due to limited interaction types, challenges in collecting and specifying context information, and difficulty in acquiring matching sound assets. We present SonifyAR, an LLM-based AR sound authoring system that generates context-aware sound effects for AR experiences. SonifyAR expands the current design space of AR sound and implements a Programming by Demonstration (PbD) pipeline to automatically collect contextual information of AR events, including virtual content semantics and real world context. This context information is then processed by a large language model to acquire sound effects with Recommendation, Retrieval, Generation, and Transfer methods. To evaluate the usability and performance of our system, we conducted a user study with eight participants and created five example applications, including an AR-based science experiment, an improving case for AR headset safety, and an assisting example for low vision AR users.
Mixed Reality, Sound, Augmented Reality, Authoring Tool
$c$s: Human-centered computing Interaction designccs: Human-centered computing Interactive systems and tools
Figure 1. SonifyAR is an AR sound authoring pipeline that generates context-matching sounds for AR events in situ. With SonifyAR, the user can first interact with the AR Scene, while SonifyAR identifies potential sound-producing events and collects context information as text. This context text, including the real-world environment information like surface material, is then processed by LLM, which guides the sound acquisition process using multiple models and methods. The acquired sounds are then used to sonify AR events.
[…]
SonifyAR overview
1. Introduction
Sound is a critical but often overlooked element in Augmented Reality (AR). AR-based sound can improve immersion and overall user experience (roginska2017immersive, ) in applications such as depth perception and task completion (zhou_role_2007, ), search and navigation (ruminski2015experimental, ), and even assistance for people with low vision  (ribeiro_auditory_2012, ). Despite its importance, current AR authoring platforms like Reality Composer (realitycomposer, ), Adobe Aero (adobeaero, ), and Unity Mars (UnityMARS2023, ) provide only rudimentary sound support. Specifically, we identified three critical gaps with existing AR authoring tools:
  1. (1)
    Limited real-world context. Existing systems typically support action triggers linked to virtual objects in AR but lack support for real-world contextual information (e.g., a virtual toy robot traversing diverse indoor surfaces like wood, carpet, or glass.).
  2. (2)
    Limited interaction specification. Existing systems provide only pre-defined interaction triggers like “Tap” and “Proximity Enter”. This limits a creator’s ability to specify interactions outside the provided options, especially those that involve environmental context (e.g., the user “slides” virtual chalk on a real-world blackboard).
  3. (3)
    Limited sound sources. Existing systems are limited by the sound assets available in their libraries and the scarcity of suitable sound resources online. Thus, AR authors struggle to find appropriate sounds for distinct AR events (e.g., reproducing the wing flutter of a virtual dragonfly or simulating the eating sound of a virtual dinosaur).
To address these challenges, we present SonifyAR: a context-aware AR sound authoring system using GPT-4 (openai2024gpt4, ) and a text2audio diffusion model called AudioLDM (liu2023audioldm, ). SonifyAR makes the following key technical advancements:
First, context collection using PbD. SonifyAR adopts a Programming by Demonstration (PbD) (10.1145/291080.291104, ) pipeline to simplify the specification of complex AR interactions. The PbD pipeline enables users to demonstrate AR sound interactions while the system automatically detects the action and collects context information. For example, if a creator wants to sonify the stomping of a walking robot, they can position the virtual robot on the target (physical) surface. As the robot walks, the collision between robot’s feet and the surface, as well as the context information like the robot’s attributes and the surface’s material, is captured by SonifyAR .
Second, context as text. To utilize AR context information stemming from multiple sources (e.g. user action, virtual object, real-world environment) and in different formats (e.g. categorical, 3D shape, image), we use text as the universal medium to encompass all context information. By describing all context with text and fitting different parts into a template (”This event is [Event Type], caused by [Source]. This event casts on [Target Object]. [Additional Information on Involved Entities]”), we represent context information of AR interactions as text snippets.
Finally, LLM-based sound acquisition. We integrate a suite of four sound acquisition methods: recommend, retrieve, generate, and transfer, all controlled by the underlying LLM. For each AR sound interaction, the context information as text is fed to the LLM for processing. The LLM then provides text prompts that control the suite of four sound acquisition methods, which automatically provide matching sound assets for the AR interaction.
To author AR sound with SonifyAR, the creator initiates AR interactions and the system automatically lists sound options based on context. For example, imagine trying to make a virtual tea cup chime crisply when being slid on a table (Figure 1). In traditional approaches, the creator would need to manually specify this sliding action and find a sound asset to match the chiming ceramic cup. But with SonifyAR, the creator could demonstrate this sliding action and the chime sounds are generated automatically for the creator to choose from.
To explore the potential of SonifyAR, we conducted two evaluations: first, a qualitative user study with eight designers to examine the usability of the SonifyAR prototype; second, we apply SonifyAR across five key user scenarios: from AR education to improving AR headset safety. Our user study highlights usability of the authoring interface and helpfulness of the generated sounds while identifying space of improvements in terms of generated sound quality. Our application scenarios demonstrate the breadth and potential of SonifyAR incorporating automatically acquired sounds seamlessly into various AR experiences—a practice that would otherwise take significant time and effort.
Our work makes three primary contributions: (1) a design space for AR sound authoring tools, which highlights existing gaps in the literature; (2) a novel context-aware AR sound authoring pipeline using generative AI, called SonifyAR, that incorporates sensed environmental cues like surface material, virtual object semantics, and user action; (3) findings from a user study with eight designers and five application examples that demonstrate SonifyAR’s potential and key advancements in this space. To our knowledge, we are the first system to offer automatic, in-context sound authoring for AR.
2. Related Work
We cover prior work in AR sound authoring, AI-based sound generation techniques, and context-awareness in AR.
2.1. AR Sound Authoring
Sound in AR provides many benefits, from directing the user’s attention and enhancing immersion (roginska2017immersive, ; zhou_role_2007, ) to creating interactive time-varying experiences and increasing accessibility (serafin_sonic_2018, ; ruminski2015experimental, ). However, as mentioned, the current landscape of AR authoring tools and platforms reveals a deficiency of support in AR sound authorship.
The existing AR authoring tools show diversity in their authoring capability (monteiro2023teachable, ; nebeling_trouble_2018, ). Some are easy to use, require no coding skills, and can only attach user-provided sound with certain AR event triggers (haloar, ; arvid, ; adobeaero, ; realitycomposer, ). Some are heavily code-based and require significant technical skills, but provide substantial flexibility in designing AR sound interactions, including the ability to set parameters like sound decay and code-specific conditions for activating sound assets(unity-website, ; unreal-engine-website, ). One common design in these tools is a Programming by Specification (PbS) methodology (monteiro2023teachable, ), where creators define AR sound content and play conditions during the design phase. This allows creators to precisely define the conditions for triggers and the consequent sound effects.
As mentioned in the Introduction, we identified critical gaps with existing AR authoring tools. These gaps are partially due to the limitation of the PbS methodology, since the triggers specified in the design stage naturally lack richness in real-world context, which is crucial for AR sound to be realistic. Also, the trade-offs between supported trigger types and the tool’s usability make most tools lack coverage for many sound-producing AR events. Additionally, all sound assets need to be manually selected or created, tasking creators with sourcing suitable sound effects to match the specific triggers. The SonifyAR system aims to address these gaps.
2.2. Sound Acquisition
Although no work has explored sound generation with AR events as input, sound generation conditioned from other input modalities like images, 3D models, text, and videos has already been widely applied. Various cross-modal generative models like RNN (Zhou_2018_CVPR, ), GAN (engel2019gansynth, ; kumar2019melgan, ; ghose2022foleygan, ), VAE (dhariwal2020jukebox, ), and most recently diffusion models (kong2020diffwave, ; liu2023audioldm, ; yang2023diffsound, ), are utilized to generate sound. Despite different model architecture, one common training goal is the mapping between latent representations of input modalities and output sound, thus sounds can be generated from prompts like images (10096023, ), videos (Zhou_2018_CVPR, ; ghose2022foleygan, ), and text (liu2023audioldm, ; yang2023diffsound, ). For example, FoleyGAN (ghose2022foleygan, ) conditions action sequences of input videos leading to the generation of visually aligned realistic soundtracks. AudioLDM (liu2023audioldm, ) utilize CLAP (wu2023large, ) based latent space to embed input text and generate matching sound with diffusion model. Unlike the cross-modal generative methods, another notable line of work is physics-based sound synthesis (raghuvanshi_interactive_2006, ; roodaki_sonifeye_2017, ; ren_example-guided_2013, ; jin_deep-modal_2020, ; jin_neuralsound_2022, ; liu_sound_2021, ; diaz_rigid-body_2022, ). Traditionally using methods like spring-mass model (raghuvanshi_interactive_2006, ) and recently improved with deep learning (jin_deep-modal_2020, ), these synthesis methods can provide realistic sound simulation given 3D model and material properties.
Besides synthesizing, retrieval methods can also acquire matching sound for input conditions. For example, Oncescu et al. (Oncescu21a, ) trained cross-modal embedding models to support text-to-audio retrieval. Multi-modality models like VAST  (chen2023vast, ) also supports text-to-audio retrieval that could potentially be used to provide sound assets for AR experiences. A recent work called Soundify (lin2023soundify, ) utilizes retrieved sound assets to sonify videos. They claim that although sound retrieval do not ensure coverage for all input descriptions, the retrieved sound assets usually surpass synthesis results in terms of quality. In this case, combining generation and retrieval in the sound acquisition pipeline seems to be a natural method that integrates merits of both. We are going to explore this possibility in this work.
Unlike other existing sound acquisition methods, AR poses unique challenges in the matching of input condition and output sound. The multi-modal nature of AR interactions, including the information of the virtual object (e.g. a toy robot with a walking animation), real environment (e.g. a living room space with a wooden table in the middle), and user intent (e.g. user taps on a virtual model), need to be carefully processed into the sound acquisition pipeline to result in matching sound assets.
2.3. Context-awareness in AR
In AR, context-awareness is an important property that allows AR content to be dynamically adjusted according to context information, such as location, scene semantics, and human factors. Context-awareness ensures that the application adapts not only to the user but also to their environment and specific needs, making the application more user-friendly and immersive (krings_development_2020, ). Previous research has explored various aspects of context-awareness in AR, highlighting its importance across different applications.
For example, the popular AR game Pokemón Go shows game contents based on user’s physical location. A more common practice in AR research is adapting virtual contents to scene semantics—essentially understanding the objects and their locations in the user’s environment. Qian et al. (qian2022scalar, ) introduced an authoring system that helps user create adaptive AR experiences to surrounding scene semantics. Lang et al. (lang2019virtual, ) analyzes scene semantics to guide the positioning of virtual agents. Similarly, Liang et al. (liang2021scene, ) places virtual pets into AR based on scene geometry and semantics. This realm of similar work (han2020live, ; tahara_retargetable_2020, ) usually captures and analyzes surrounding environment, and designs algorithms that determines possible interaction between virtual content and real-world scene. Lindbauer et al. (lindlbauer2019context, ) push the AR context-awareness further by including task and cognitive load as context to adjust level of detail of AR virtual interface.
Although there has been substantial discussion in the realm of AR context-awareness, we observed that prior work generally focuses on the visual modality of AR, neglecting sound. Additionally, existing work primarily concentrates on adjusting and arranging virtual content rather than generating new content. We aim to fill this gap by designing a sound generation pipeline that processes context information to produce matching sound assets.
3. Design of SonifyAR
To design an effective sound authoring system that generates matching sounds for AR interaction, three main research questions need to be answered.
  • What are the target AR interactions that should produce sound effects?
  • How can we efficiently acquire sound assets for these AR interactions?
  • How can these AR interactions be specified and represented in the authoring process?
We aim to elaborate on these questions and the respective design rationale in this section.
3.1. The Triad of User, Virtuality, and Reality
We began by examining the design space of AR interaction sounds. Drawing on Jain et al.’s (jain_taxonomy_2021, ) sound taxonomy in VR, we attempted a similar analysis for AR sound interaction. We first investigated several widely used code-free AR authoring tools for their AR sound capability, including Adobe Aero (adobeaero, ), Apple Reality Composer (realitycomposer, ), and Halo AR (haloar, ). We found that these tools support AR sound related to either virtual content interaction or AR-based spatial anchors, e.g. play sound when users get close to a certain position, scan a QR code, or tap on a virtual model.
The research literature in AR sound (dam2024taxonomy, ; rakkolainen2021technologies, ; filus2012using, ), however, envisions a much broader scope. E.g. Rakkolainen et al. (rakkolainen2021technologies, ) have classified AR audio into five classes: ambient, directional, musical, speech, and noise. Dam et al.’s (dam2024taxonomy, ) taxonomy of Audio Augmented Reality (AAR), defines AR-based sound across Environment Connected (sound maintaining awareness and interaction with physical environment), Goal Directed (sound assisting users’ primary goal), and Context Adapted (sound adaptive to users’ immediate reality), as well as their intersection. This taxonomy emphasizes that AR sound is often rooted in contextual information from the real world environment and user intention.
Drawing on these prior work, we introduce a novel AR sound design space that helps emphasize the sonic interplay between three key factors of AR experiences: reality, denoting the physical environment in the real world that hosts the AR experience; virtuality, referring to the virtual object(s) placed in the AR experience; and the user, representing the individual interacting with the AR experience. Below, we expand on each element of this proposed design space and indicate current levels of support in the existing code-free AR sound authoring tools (adobeaero, ; realitycomposer, ; haloar, ).
Figure 2. The sound-producing opportunities in the triad of User, Virtuality and Reality.
Virtuality (partially supported): Virtuality refers to sounds that accompany AR events involving only virtual objects. This could be bound to an object’s change of status (e.g. a notifying sound when a virtual object shows up), from an object’s animated behavior (e.g. the mechanical noise made by a virtual dinosaur roaring), or the interaction between multiple virtual objects (e.g. two virtual balls clacking when collide). Current authoring tools support status change and animation playing as sound initiating triggers, but do not support sound for interactions between virtual objects.
User-Virtuality (well supported): User-Virtuality are the corresponding sounds that react to the users’ actions with virtual objects, e.g. a virtual dog barks when users tap on or get close to it. In all of our investigated AR creation tools, these user-initiated triggers like “Tap” and “Proximity Enter” are well supported.
User-Reality (not well supported): User-Reality refers to sounds that accompany user interactions with the physical environment via their AR devices. These sound feedbacks can improve users’ understanding of the surrounding environment. E.g. a tap sound when users tapping on a real-world table surface on their phone screen. None of our investigated AR authoring tools support user-reality actions as triggers.
Virtuality-Reality (not well supported): Virtuality-Reality denotes sound feedback that plays crucial roles in enhancing realism when virtual objects interact with the real-world environment, e.g. the crisp stomping sound of a virtual robot when it walks on a real-world glass surface. Although physics simulation between virtual models and real-world surfaces have been supported, their sound feedback remains unexplored in AR authoring tools.
User-Virtuality-Reality (not well supported): User-Virtuality-Reality are the sounds that accompany user actions involving both virtual and real elements. For example, the material-aware sliding sounds when users apply a virtual scraper on different real-world surfaces (e.g., wooden table, painted wall, or glass window). None of our investigated authoring tools can support the specification of this complex interaction.
Others Note that we excluded the domains of Reality and User from the enumeration since our interest lies in events where the AR system (e.g. smartphone or AR headset) generate the sound. Thus, sounds naturally generated by user and reality are excluded from our discussion.
Based on the above analysis and as illustrated in Fig. 2, existing AR authoring tools fail to offer functional AR sound authoring support for many of the identified dimensions, especially those involving Reality. SonifyAR aims to cover this gap.
3.2. Acquisition of AR Sound
While the above design space helps highlight the dimensions of AR sound and current authoring support, below we reflect on key challenges related to the limitations of sound acquisition methods in terms of handling these different dimensions.
For example, a mechanical noise of a virtual robot (virtuality) can be easily sourced from local or online sound databases. However, more specific sounds, like a virtual steel ball hitting a physical concrete wall surface or a virtual racecar jumping into a backyard pool (virtuality-reality) requires physical world understanding and material awareness. The space of possible sound effects here is near infinite. Thus, in such cases, generating sounds using text-to-sound models is preferable.
Conversely, certain highly recognizable sounds, like a dog barking, pose difficulties for current generative models, often resulting in outputs that are unsatisfactory and noisy. The complexity in user needs and trade-offs in different methods suggests that the combination of both generation and retrieval could serve as a practical solution.
3.3. Event Representation
To combine generation and retrieval methods, one challenge is how to condition this suite of different sound acquisition methods. While visual formats such as images (10096023, ) and videos (Zhou_2018_CVPR, ; ghose2022foleygan, ) have been widely used as inputs for sound generation models, they fall short in fully capturing AR context information, particularly user actions. To overcome this, we explore the use of text as a universal representation for sound acquisition inputs, which has the following benefits. First, text can sufficiently and precisely convey necessary context information and the specifics of sound-generating events (e.g.,“User slides a ceramic cup on a wooden surface”). Second, given that the AR experience operates within a hardware-software system with multi-modal sensors (e.g., camera, IMU, GPS), we can leverage this system to monitor and log events within the AR experience and summarize them into descriptive text. Third, recent advancements, as demonstrated by works like HuggingGPT (shen2023hugginggpt, ) which uses a LLM to control AI tasks, showcase the practicality of using LLMs as controllers to process text for sound generation.
4. SonifyAR Implementation
Figure 3. Overview of the pipeline of SonifyAR. Our system monitors and logs context information of AR events, which includes the event type, the subjects and objects (virtual or real-world), and the attributes of the involved elements like their materials. This information is compiled into a text template and then processed by our LLM controller to acquire sound assets. The results are subsequently presented in our selection panel.
We introduce SonifyAR, a novel PbD sound authoring framework that uses context recognition and generative AI to create personalized, context-sensitive sounds for AR interactions (Figure 3). SonifyAR consists of three primary components : (1) Event Textualization: Every user action is recognized as an AR event, and the context of these events is transformed into textual descriptions. (2) Sound Acquisition: An LLM-controlled sound acquisition process that utilize four methods to produces sound effects based on the context. (3) User Interface: An interface that allows users to experience the AR scene and provides the capability to view, modify, and test sound effects for AR events.
4.1. Event Textualization
In SonifyAR, we utilize a PbD authoring framework to capture potential sound-producing AR events and compile the context information into text. When users interact with the AR space, SonifyAR captures AR events that can lead to sound feedback. Here we define an AR event as an user action and its subsequent result (e.g. when a user taps a virtual object to trigger its animation). For each AR event, we describe it with an event type (e.g. tapping an object ), action source (e.g. user or virtual object), and action target (e.g. virtual object or real-world plane). Information about the involved parties, such as virtual objects or real-world planes, will also be included in the context information. This collected event data is then transformed into text with a template.
Figure 4. SonifyAR’s event textualization. Left: a user tapping on a virtual cup through the SonifyAR interface; Right: the textual context information extracted by SonifyAR’s internal PbD framework.
Event Types.
To support a broad range of AR sound interaction, we implemented six event types:
  1. (1)
    Tap Real World Structure: A user taps on a real-world structure through the AR interface.
  2. (2)
    Slide: A user holds the virtual object and slides it on a real-world surface.
  3. (3)
    Collide: A virtual object collides with another or with a real-world surface. We implemented a specialized collision detection mechanism as introduced in subsection 4.4
  4. (4)
    Show Up: A virtual object shows up in the AR experience.
  5. (5)
    Tap Virtual Objects: A user taps on a virtual object through the AR interface.
  6. (6)
    Play Animation: A virtual object plays an animation.
Scene Context Understanding.
SonifyAR utilizes ARKit (apple-arkit-doc, )’s plane detection functionality (apple-arkit-plane-tracking, ) to assess the surrounding environment. Since most AR experiences are anchored to planes, the detected plane information serves as a crucial context. To enrich such information, we employ the Deep Material Segmentation model (upchurch_dense_2022, ) to segment the scene and identify the material of planes (Figure 3), which helps produce realistic sound effects when an AR event involves this plane. Currently, we support six materials: wood, carpet, concrete, paper, metal, and glass. Surfaces that do not fall into one of these six categories are labeled as unknown surface.
Virtual Object Understanding
The semantics of the virtual object is also crucial for sound generation. To ensure sound assets align with the virtual object’s material and its state, we collect a text description for all the virtual objects (e.g. ”This model is a toy robot made of metal.”) and their animations (e.g. ”A toy robot walks.”). These descriptions can be provided by the asset creator or, if necessary, we prompt the user to add relevant details. Any AR event involving the virtual object or its animation will incorporate these text descriptions.
Event-to-text
To aggregate the multi-source context information described above, we employ a simple text template: ”This event is [Event Type], caused by [Source]. This event casts on [Target Object]. [Additional Information on Involved Entities]”.
In our template, Event Type is the type of event described by the type name, as listed above. Source refers to the subject of the event, which could be the user when the event is directly triggered by user, or virtual object when it interacts with real-world environment. Target Object is the object of the event, e.g. the plane that gets tapped on, or the animation that gets played. Additional Information on Involved Entities include details that further elucidate the source object, the triggered event, and the target object. Examples include material descriptions and animation details. We leave the corresponding entity field blank when events are missing a target object or additional information.
During user interactions, the system actively monitors and logs the events happening in AR space. The context information is fetched and plugged into the text template, crafting a coherent description that reflects the user’s interaction within the AR environment.
4.2. Sound Acquisition
Figure 5. SonifyAR’s sound acquisition pipeline.
SonifyAR utilizes the LLM to automatically retrieve or generate context-matching sound assets of an AR event. Inspired by HuggingGPT (shen2023hugginggpt, ) and Visual ChatGPT (wu2023visual, ), we utilize the LLM as a controller of multiple sound authoring methods. The LLM takes the text description of the event as input and replies with commands for multiple sound acquisition methods (Figure 5). At our current stage, we support four major sound authoring methods: local recommendation, online retrieval, text2sound generation, and text-guided sound transfer.
Local Recommendation
The LLM can recommend sound assets stored in the local database based on semantics in the event description. Similar to other AR authoring tools, we collect a set of sound effects from Adobe Audition sound effects (adobe-audition, ), each labeled with a descriptive file name, like “Crash Aluminum Tray Bang” or “Liquid Mud Suction”. The entire list of sound file name is provided to LLM. When provided with the event context, the LLM recommends the top five most suitable local sound effects based on their file names. It then returns the selected sound effect in the format of method1recommend:FILENAME, where the FILENAME would be replaced with the actual filename. Upon receiving, SonifyAR parses the filename and adds the top corresponding sound as one of the sound options for the event. User can also long hold such option to reveal other results from the top5 recommendation.
Online Retrieval
We also expand our retrieval capability with an online sound asset database called FreeSound (freesound, ). We use the FreeSound API, which returns a list of best matching sound effects based on a given query. The queries are condensed versions of full event descriptions generated by the LLM, returned in the format method2retrieval:PROMPT, with PROMPT replaced by the specific search query. The returned results are JSON strings containing information on all search results. We select the top five matches from the entire set of sound effects returned by the API, which are then downloaded and presented to the user as sound options.
Sound Generation
We adopt a audio diffusion model AudioLDM (liu2023audioldm, ) to generate sound effects with text prompt. We ask the LLM to compress the event text description into a shortened generation prompt (returned in the format method3generation:PROMPT, where the PROMPT would be replaced with the generation prompt). Upon receiving such a command from the LLM, SonifyAR sends the prompt to the AudioLDM model, requesting text2sound generation. These newly generated sounds are then presented to user as sound options.
Sound Transfer
Besides text2audio generation, AudioLDM is also capable of performing text-based sound style transfer. Specifically, for events like tapping, sliding, or colliding, instead of generating a new sound from scratch, SonifyAR uses a default sound effect and initiate a style transfer operation with a text prompt provided by the LLM. This approach allows the output sounds to match the length and rhythm of the input sounds, so that they can be well-timed with actions. Furthermore, when users wish to further modify any provided sound assets, SonifyAR offers text-based style transfer as a fine-tuning and customizing option.
All sound authoring processes listed above operate concurrently in the backend. This ensures uninterrupted AR experience.
4.3. User Interface
Figure 6. SonifyAR’s authoring interface. Left: SonifyAR’s phone-based AR interface. Right: SonifyAR’s authoring panel.
As a PbD authoring framework, the SonifyAR application initiate directly with an interactive AR experience (Figure 6 left). Users can freely interact with the scene, performing actions like moving virtual objects or interacting with the real-world surfaces. When an AR event is detected, a text label appears on the top-left screen to inform the user (Figure 6. A) and the sound acquisition component is activated to generate candidate sound effects for the detected events.
By clicking a button (Figure 6. D), users can prompt an editing interface (Figure 6. F) that overlays on the AR scene. The interface includes all unique AR events detected (Figure 6. G) and the automatically generated sound assets (Figure 6. H), enabling users to explore the detected events and their corresponding sound effects. Users can click on a sound effect to preview it and double click to select and activate it during the AR session. After confirming their choices, users can hide the authoring panel to resume the AR experience and test the selected sounds. Users can always access the editing interface to modify their choices.
Furthermore, users can long press a sound asset to call out a menu with a suite of exploratory options. For recommended or retrieved sounds, the menu includes an option to list all other sound assets in the top five recommendations. For all sound assets, the menu provides a “style transfer” and a “generate similar sounds” feature, enabling users to style transfer audio effects or to generate similar sounds based on the selected sound. When this feature is selected, users can type a simple text prompt to guide the sound generation process. This feature enables users to iteratively refine sound effects and explore new sound variations.
4.4. Technical Implementation
We developed SonifyAR as a mobile app with a user interface that enables AR interaction and sound authoring actions. We also implemented back-end servers for scene understanding and sound generation. The current SonifyAR demo is built with Apple’s ARKit with Swift 5.7 and Xcode 14, and runs as an iOS application on iPhones. Due to the enhanced performance of plane detection on devices with LiDAR (available in iPhone’s Pro lineup starting from iPhone Pro 12), we developed and tested the app on an iPhone 13 Pro Max running iOS 16. We implemented the collision detection between virtual objects and planes with ARKit’s physics simulation (apple-scenekit-physics-simulation, ) function. To enable collision detection for animated virtual object parts (e.g. the stomping of a robot foot on the plane), we bind colliders to the joints of the virtual object’s animated bones.
We use GPT-4 (openai2024gpt4, ) as the LLM in our system. We adopt the DMS model (upchurch_dense_2022, ) for material segmentation because of its detailed material labeling, fast processing time, and high accuracy. For sound generation, we use the AudioLDM model (liu2023audioldm, ) due to its state-of-the-art performance and its versatility in text2sound generation and text-guided sound2sound transfer. All models are hosted on a server equipped with an NVIDIA RTX 4080 GPU and CUDA 12.2.
5. User Study
To examine the usability and authoring performance of the SonifyAR system, we conducted a usability study across eight participants, who used the SonifyAR application in a fixed experimental setting and provided ratings for the system.
5.1. Participants
We recruit eight participants (six male and two female, aged between 26 and 33) with varying experience in AR. We inquired about their prior experience in AR before the study. Out of the eight participants, six had prior AR experience, and three had experience in AR authoring tools.
5.2. Procedure
The usability study was conducted in a small meeting room and comprised four steps. We first introduce our study goal with the participants and collect answers to our demographic questions. We then demonstrated our SonifyAR app with an example AR experience. During this phase, participants had the opportunity to ask any questions regarding SonifyAR’s functionalities. Afterwards, participants were asked to independently explore the SonifyAR app. We provided a walking robot model which can be added to indoor surfaces. This model starts walking when tapped. We asked the participants to explore this AR asset freely and use our automatic sound authoring pipeline to create sound effects with the model until they are satisfied with the results. The entire usage process was screen-recorded and the user operations were logged. The final step involved getting participants’ feedback regarding the usability, helpfulness and technical performance of the SonifyAR system based their personal experiences. Participants also provided reasoning and insights to support their answers. The questions and rating results are shown in Figure 7.
5.3. Results
All participants were able to complete the AR sound authoring task using SonifyAR without difficulty. On average, the authoring process took 406 seconds (SD=137s) to finish. Participants tested an average of 56 (SD=10.48) sound assets, with an average of 6.4 sounds (SD=1.67) assigned to the authoring results.
Feedback on the SonifyAR app was largely positive. All eight participants expressed favorable impressions of the tool. They highly agreed that the SonifyAR tool would be helpful to AR authoring process (Q1,AVG=6 out of 7, SD=0.93). There was also a general willingness for using SonifyAR in their own AR creation practice (Q2, AVG=6.25, SD=1.16). Participants agreed that the sound interaction involving real-world surfaces will improve the immersiveness of AR experiences (Q3, AVG=6.38, SD=1.06). Additionally, they preferred the automated sound authoring process over the manual search for sound assets (Q5, AVG=6.13, SD=0.99). The only area for improvement was the quality of the generated sound (Q4, AVG=4.75, SD=1.39). However, this could be enhanced with the introduction of a more advanced sound generative model, which can be easily integrated into our system. The full list of questions and the respective Likert-type results can be viewed in Figure 7.
Figure 7. Likert-type questions and results of our usability study
6. Applications
To demonstrate the potential and technical capability of SonifyAR, we explore five example application scenarios that utilize SonifyAR’s sound authoring pipeline in AR applications. All examples are also shown in supplementary videos for better presentation.
6.1. Education
AR have been widely explored in STEM education (sirakaya2022augmented, ). As existing educational applications try to simulate and visualize physical or chemical phenomenon in AR (radu2019can, ; irwansyah2018augmented, ), we envision that SonifyAR can help add sound into the simulation and improve the blending between real environment and virtual contents.
To showcase such possibility, we implement an AR physics experiment with SonifyAR demo application. We create a 3D model to show the conservation of momentum, which is a slope and two metal balls located at top and bottom of the slope (Figure 8). Upon initiating the experiment, the top ball rolls down the slope and collides with the bottom ball. This collision causes the bottom ball to move forward and eventually fall, while the top ball comes to a stop. With SonifyAR, collision events are automatically detected, and material-specific sound effects, such as the clashing of two metal balls or a metal ball dropping onto table, are seamlessly generated. The resulting sound effects can be experienced in the supplementary video.
Figure 8. An AR physics experiment sonified by SonifyAR. (A) The slope and metal ball model and simulation process. (B), (C) Context information of two collisions in the simulation. (D) The acquired sound assets for the collisions.
6.2. Accessibility
AAR (Audio Augmented Reality) has been explored as accessibility assistance for blind or low vision people (ribeiro2012auditory, ; coughlan2017ar4vi, ). By sonifying real world environments and virtual contents, blind or low vision (BLV) people can better interpret visual information in both reality and virtuality. Following this thread of work, we envision SonifyAR being able to assist on accessibility in AR experiences in three ways. First, by supporting AR sound authoring, SonifyAR encourages creators to add sound effects in AR experiences, which can help BLV users consume visual contents. Second, by providing context-aware AR sound in 3D audio, people with low-vision can better navigate virtual objects in AR environment. Third, by enabling user to interact with real-world surfaces via AR (e.g. tapping on a real world surface), user can explore the surrounding space via the AR interface.
Figure 9. Robot walking on different surface with different sound. (A), (B) Robot walking on a wooden surface, and the corresponding context information. (C), (D) Robot walking on carpet surface, and the corresponding context information. (E), (F) The acquired sound assets.
We use a toy robot 3D model and the SonifyAR demo app to showcase the accessibility benefits. The toy robot (Figure 9) can be placed on any indoor surface and has a walking animation. By authoring different bouncing sound effects for different surfaces (e.g. wood and carpet), user can better perceive the location of the robot while it is walking.
6.3. Sonify Existing Apps
Due to the extensibility of the text template used in SonifyAR’s sound acquisition pipeline, we foresee its application in existing AR apps if integrated as an extension. If implemented at the SDK level (e.g. ARKit and ARCore), SonifyAR can automatically capture textual description of AR events and provide sonification using the automatic sound acquisition results.
Figure 10. AR dinosaur sonified by SonifyAR. (A) AR Camera app interface with a virtual dinosaur playing eating animation in the scene. (B) The context information of the scene. (C) Text prompt and sound output from SonifyAR’s pipeline.
We explore this application possibility with a Wizard of Oz (WoZ) prototype on a phone-based AR application called ARVid (arvid, ), which is an AR app that puts 3D animated objects into real world environment. We select a dinosaur animation as an example (Figure 10), and conceptualize that SonifyAR pipeline is adopted into the AR infrastructure. By manually feeding the context information (e.g. “The dinosaur is eating”) into the SonifyAR pipeline, we generated sound assets that well-match the 3D animations. The results are shown in the supplementary video.
6.4. Sonify virtuality-reality interaction on AR/MR headset
As recent AR/MR headsets become increasingly popular in entertainment and productivity tasks, manufacturers have established safety guidelines, such as setting up safety zone and clear up indoor spaces. We envision SonifyAR contributing to this topic by creating material-aware sound effects during headset usage when application windows or virtual object intersect with real world surfaces or objects.
Figure 11. Sonification of Vision pro’s window operation. (A) User place a virtual window on sofa. (B) Context information of this action. (C) Text prompt and sound output from SonifyAR’s pipeline.
We implement a WoZ prototype based on Apple Vision Pro’s user interaction (Figure 11). When using Vision Pro, user can actively adjust the position of application windows by hand gestures. SonifyAR can enhance this interaction by generating AR sounds when virtual windows are placed on or intersect with indoor surfaces like walls, desks, and floors. By processing these interactions and applying sound transfers with text prompts such as “a virtual window bumps into a wall” and “a virtual window collides with a carpet”, SonifyAR aims to improve the permanence of virtual windows and increase awareness of surrounding objects as users navigate through virtual windows. See supplementary video for more direct inspection of the sound effects.
6.5. Sound asset generation for traditional AR authoring process
Besides serving as a PbD authoring experience, SonifyAR’s sound authoring pipeline also have the potential of improving existing PbS AR authoring tools in sound assets provision. As the mainstream AR authoring tools like Adobe Aero and Apple Reality Composer already represent their AR interaction as text like “Tap & Play Sound” and “Proximity & Jiggle”, these specifications can be fed into SonifyAR’s text-based sound authoring pipeline. In this case, SonifyAR works as a plugin of these tools and provide matching sound assets for AR interactions.
Figure 12. WoZ example of SonifyAR being applied in Reality Composer’s authoring process. (A) A machine model in Reality Composer’s authoring interface. (B) Reality Composer’s interface for AR sound. (C) Context information of the event in B. (D), (E) Two sound assets generated and retrieved with SonifyAR’s pipeline. (F) The acquired sound being used as recommended options.
We implement a WoZ prototype using Apple’s Reality Composer and a 3D model of an animated machine (Figure 12). When authoring the “Play Sound” behavior of this model, context information like the model description, triggering condition, and animation description, can be fed into SonifyAR’s text-based authoring pipeline and lead to two sound assets from generation and retrieval methods. These two sound assets will appear at top of the sound asset selection as “SonifyAR Recommended”. See supplementary video for more details.
7. Discussion and Future Work
In this paper, we designed and implemented an AR sound authoring pipeline that automatically generates sound assets based on context information. We evaluated and discovered the opportunities in the AR sound interaction space, and implemented an LLM+Generative AI pipeline for sound acquisition. We tested the pipeline with a usability study on our prototype and further discussed the application opportunity with five examples.
7.1. Limitations and failure cases
While SonifyAR was positively evaluated in our user study and demonstrated broad application potential, we observed limitations in the system adaptability and encountered failure cases in multiple aspects.
Firstly, since the sound acquisition pipeline relies on LLM as the context collector and overall controller, errors of the LLM can lead to failure cases in sound acquisition. For example, the LLM can extract the wrong keywords in the context information and prompt sound generation model with inaccruate text input (e.g. LLM provide prompt of “sliding plastic on carpet” while the sliding model is actually a metal toy robot) that lead to non-matching outputs.
Secondly, although AudioLDM (liu2023audioldm, ) is the state-of-the-art text2audio generator and performs decently in most cases, it can sometimes produce subpar outputs that are noisy or do not match prompts. Based on our experiences, such inconsistency in performance can be hard to predict thus cannot be completely solved by prompt engineering.
Thirdly, some AR interactions can be highly time-sensitive (e.g. objects colliding, the playing of animations that shows very clear action sequences) and require sound assets being closely synchronized with visual contents. Currently, we provide example sound assets for AR interactions of “Colliding”, “Tapping” and “’Sliding’. By conducting sound style transfer using well-timed example sounds as inputs, the output sounds can be formatted to be in sync with the AR interaction. However, for animations, we currently do not have methods to sync the generated sound with the key frames in animation, so some generated sounds can be off-sync with animation. We believe this 3D animation-to-audio synthesis challenge can be a standalone topic for future researchers to solve.
7.2. Expansion of framework
Currently, we implement the automated AR sound acquisition with six example AR event types, as listed in section 4. This selection is primarily for proof of concept. Future work, or other systems adopting SonifyAR’s text-based sound acquisition pipeline, could easily incorporate additional types of AR interactions, provided they have clear and concise text descriptions. For example, event types used in Adobe Aero and Reality Composer, such as “Spin”, “Emphasize”, “Expand”, or “Jump”, can be well-incorporated into the text template of SonifyAR. The flexibility and extensibility of our text-based pipeline is also explored in our application examples.
7.3. Future work
We plan to address the following improvements for future work.
First, our current SonifyAR app is a simplified tool that exclusively supports the authoring of AR sound and uses pre-crafted AR scenes as input. For wider applicability that reflects the real-world needs of AR content creation, we recognize the need to incorporate the AR sound authoring pipeline into the entire AR content creation process. Our future goals include integrating the SonifyAR framework with established tools like Reality Composer and Adobe Aero.
Secondly, SonifyAR adopts a straightforward process of using a text template to compile context information collected from multiple sources, including virtual object semantics and real world semantics. Currently, we use one material segmentation model to perceive real-world context information. In future developments, we aspire to incorporate more sophisticated computer vision models, or multi-modal foundational models, for deeper understanding of both virtual objects and real world environment, enhancing the context acquisition process with greater precision and broader adaptability.
Lastly, the current sound generation output of SonifyAR presents opportunities for further refinement. Recognizing the rapid advancements in the AI research domain, our system has been designed as a modular framework, ensuring that as more advanced models emerge, they can be seamlessly integrated. Every model (e.g., LLM, CV, Sound Generation) within SonifyAR is replaceable, allowing SonifyAR to easily adapt and stay at the forefront of innovation in this field.
8. Conclusion
In this paper, we present SonifyAR, a context-aware sound authoring system designed for sonifying AR events. Our work implements a PbD authoring pipeline which enables sound asset acquisition using context information. With SonifyAR, users could demonstrate AR interactions and have AR sound assets automatically acquired. Our studies validate the usability and overall performance of our system. Our five application examples further widens our application potential. We believe SonifyAR can be applied in application fields like education, BLV assistance, and AR headset safety. We also envision it incorporated into existing AR authoring tools to strengthen their sound authoring capabilities. This research advances literature in AR sound authoring, while also opening up new research avenues for the application of LLM and generative models in AR systems.